OpenStudy (ajtran):

Find the equation of the plane containing the points (2,0,0) (0,6,0) and (0,0,5) in the form of ax+by+cz=D

OpenStudy (adrimit):

what is a vector normal to the plane passing through the given points?

OpenStudy (michele_laino):

hint: if we substitute the coordinates of the given points into the equation: \[ax + by + cz = d\] we get the subsequent linear system: \[\left\{ \begin{gathered} 2a = d \hfill \\ 6b = d \hfill \\ 5c = d \hfill \\ \end{gathered} \right.\] which has \(\infty\) solutions. Please solve such linear system